# Can Neural Decompilation Assist Vulnerability Prediction on Binary Code?

### v1

@@ Abstract.
Vulnerability prediction is valuable in identifying security issues more efficiently, even though it requires the source code of the target software system, which is a restrictive hypothesis. This paper presents an experimental study to predict vulnerabilities in binary code without source code or complex representations of the binary, leveraging the pivotal idea of decompiling the binary file through neural decompilation and predicting vulnerabilities through deep learning on the decompiled source code. The results outperform the state-of-the-art in both neural decompilation and vulnerability prediction, showing that it is possible to identify vulnerable programs with this approach concerning bi-class (vulnerable/non-vulnerable) and multi-class (type of vulnerability) analysis.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Software security engineeringccs: Security and privacy Vulnerability management
@@ 1. Introduction
Table 1. Intermediate steps from binary file to vulnerability prediction.
Binary ASM x86 Decompiled C++ Vulnerability
0101 1001 0111 ... <funct0>: endbr 64 push rbp ... void funct0(){ int data; data=0; ... • The function is vulnerable.
• The vulnerability is linked to CWE 190.
Neural decompilation is an emerging method that leverages modern deep learning techniques, such as transformers and self-attention, to translate binary code back into source code. One of the key benefits of this approach is that the control flow structure of the decompiled program resembles the original source code, in a closer way than disassemblers and traditional decompiler tools, such as Ghidra (Ghidra, ). It is a data-driven approach, where the neural network learns the typical patterns of source code structures. Gaining access to the source code offers several advantages, such as enabling vulnerability detection, malware recognition, re-engineering of legacy systems, and facilitating static analysis. Moreover, binary decompilation enables vulnerability prediction in “black-box” vulnerability assessments (e.g., proprietary software), where the assessor attacks the system from the same perspective of the attacker, without source code.
@@
In recent years, several solutions have emerged to predict vulnerabilities through binary code analysis, which can analyze a binary file even if its source code is unavailable (dahl2020stack, ; eschweiler2016discovre, ; grieco2016toward, ; pewny2015cross, ; schaad2023deep, ; zheng2019recurrent, ). These solutions rely on assembly code obtained from disassemblers, or on intermediate code representations such as the one from the LLVM toolchain (hasabnis2016lifting, ). However, these representations do not fully capture the semantic richness provided by the source code. The difficulties lie in several aspects: i) lack of structural information about the layout of data and code, which can be interleaved; ii) complexity and heterogeneity of the target architectures (which may have highly variable instruction sets and calling conventions); iii) optimizations performed by the compiler, which can significantly modify the program structure, (e.g., loop optimizations); and iv) lack of function and variable identifiers, which are typically stripped away from commercial software binaries (arakelyan2021bin2vec, ). Moreover, binary-oriented solutions, such as traditional decompilers, present intrinsic limitations due to the semantic gap between source and binary code (e.g., loss of the control flow structure during compilation).
In this work, we prove the potential of identifying software vulnerabilities in binary files via neural decompilation (Figure 1), without needing to access the source code of the target software. Table 1 shows the output of each step of the proposed approach. The results of this study show that:
@@
The remainder of the paper is structured as follows. In Section 2, we elaborate on background concepts about the target problem. In Section 3, we present the experimental approach for vulnerability prediction via neural decompilation. In Section 4, we illustrate the experimental analysis for the decompilation and vulnerability prediction tasks. In Section 5, we elaborate on the experimental evaluation results. In Section 6, we examine related work. In Section 7, we discuss the threats to validity. Section 8 concludes the paper.
@@ 2. Background
The main complexities of this work lie in the decompilation process, i.e., the transformation of executable code into source code, which is not completely reversible. Several decompilers are available in this field (e.g., Ghidra (Ghidra, ), RetDec (RetDec, )), achieving good levels of accuracy, however, they come with high costs in terms of knowledge of the specific language, and generation of control flow graphs. Traditional decompilation methods typically leverage an intermediate language (IL), independent from the architecture used to trace the source code, which increases the cost of developing such a system even more. In addition, the resulting decompiled files are not highly readable, as they closely reflect the structure of the assembly code.
2.1. Semantic Gap
To understand the issues involved in decompilation, we recap the steps of the compilation process, from a program written in a high-level language to the corresponding executable file. First, the program is compiled into an assembly file. Next, the assembly is transformed into a binary object file, i.e., machine code not yet executable because of unresolved function calls reported in the relocation information. Therefore, the linker must transform the object file into an executable file by linking standard library functions and other programs.
In contrast, transforming executable code into source code requires different methodologies and tools. The difficulties lie in the operations of disassembly and decompilation. Disassembly consists of lifting binary code to assembly code. There are several open-source tools available for this task. On the other hand, there is not a predefined reverse procedure for decompilation: once a program is compiled into low-level code most of the information is lost.
This issue stems from the semantic gap between high- and low-level code: machine code does not provide function-level abstraction since function calls are implemented through call labels and jumps that allow the execution of loops and conditional constructs. This asymmetry causes the loss of the control flow of the original program, which is quite challenging to re-construct from the binary.
Another problem is that compilers may rearrange code differently and transform it according to optimization rules. For example, a compiler may implement multiplication/division through a series of sums/subtractions. Also, an assembled file typically contains operating system directives that do not affect the particular program. Unaware of such complications, a decompiler would only translate what is reported in the assembly code.
2.2. Dataset
One of the complexities in dealing with the translation and vulnerability prediction tasks lies in finding an appropriate dataset. The dataset should meet the following requirements:
@@ 2.2. Dataset
In addition, we will need to pre-process the collection of programs before feeding the decompilation model. The output of such a model will be the input to the vulnerability prediction stage.
@@
ReVeal (arewethereyet, ) $18,169$ No N/A
Devign (zhou2019devign, ) $26,037$ No N/A
Juliet (Juliet, ) $128,198$ Yes Yes
BigVul (bigvul, ) $264,919$ No Only for 8,783 samples
@@
Table 2 presents an overview of the C/C++ datasets for vulnerability prediction available in the literature. It is possible to notice that datasets such as BigVul (bigvul, ) or DiverseVul (chen2023diversevul, ) provide a significant number of samples. However, they do not provide compilable functions, hence violating the first dataset requirements, i.e., compilability. On the other hand, they map the vulnerabilities on the CWEs taxonomy for small fractions of the samples. ReVeal (arewethereyet, ) and Devign (zhou2019devign, ) do not fulfill both requirements: their C/C++ code is not compilable and their vulnerabilities are not mapped on any standardized taxonomy. For these reasons, the only suitable dataset for our study is Juliet C/C++ 1.3 (Juliet, ), a synthetic collection of C/C++ programs from the Software Assurance Reference Dataset (SARD) (SARD, ), developed by NIST (NIST, ). The Juliet test suite was first released in 2010, gaining increasing popularity for benchmarking static analysis tools and vulnerability prediction tasks (zheng2019recurrent, ; schaad2023deep, ; li2021vuldeelocator, ; li2018vuldeepecker, ; jain2023code, ; gao2018comprehensive, ) since it includes both vulnerable and non-vulnerable programs. The structure of the dataset provides:
  • $64,099$ compilable test cases for 118 CWEs. For each CWE, a test case is composed of a program vulnerable to the target CWE and a non-vulnerable one.
  • A variable number of test cases for each CWE, ranging from a few units to thousands of examples.
  • A single test case may consist of a single C/C++ file or have a modular structure consisting of multiple files, with header files containing the functions’ definitions.
  • When compiled, each test case generates two executables, a bad and a good version, respectively, a version that does not handle the vulnerability and one that implements preventive measures.
@@ 3. Methodology
  • To avoid translating binary code to high-level code directly, we disassembled the binary into assembly code. Consequently, the input for the decompilation model will be assembly code, further pre-processed in the next steps.
@@
Our experimental approach revolves around five phases: Binary Generation, Disassembly, Pre-processing, Neural Decompilation, and Vulnerability Prediction, as in Figure 2.
3.1. Binary Generation
This stage consists of compiling the programs under analysis to generate two executables for each of them, a bad and a good version (JulietGithub, ). Respectively, bad executables are vulnerable to the target CWE, while good executables implement countermeasures to prevent it.
As introduced in Section 2.2, Juliet contains $64,099$ samples that cover 118 CWEs, however, a subset of such programs is not compilable due to issues with external libraries. It is worth noting that the programs were compiled without specific optimization options (JulietGithub, ), as performed by other works in the field (schaad2023deep, ). Consequently, the resulting test cases ($41,760$) cover 89 CWEs (75% of the total number) that span several categories of vulnerabilities, e.g., memory and resource management, injection flaws. Table 3 reports some statistics about the compiled dataset.
3.2. Disassembly
This stage lifts the binary code to assembly code. This transformation depends on the target architecture for executables. Several tools are available for this task: we chose objdump (objdump, ), a popular and easy-to-use tool. The format of the executables is elf64-x86-64.
3.3. Pre-processing
Assembly code. The machine code files are not yet ready for training since they are extremely long and contain library functions irrelevant to our purpose. An assembly file consists of several sections, with the code section being .text. We removed all sections except .text since we are only interested in the algorithmic logic.
At this stage, it is still necessary to remove library functions defined within the same section, operating system directives, stack initialization, and the main function, which is constant in every program, making its contribution to learning negligible. What is left are exclusively the (non-)vulnerable functions of interest. We aim to let the model learn algorithm recognition and structural patterns related to functional programming logic. We also replaced function names with standard names for each test case. This operation allows the model to focus on the function content rather than the function name.
The main benefit of this stage concerns the length of the assembly code for each test case, which is drastically reduced by more than 90% (on average). Then, the assembly code is ready for the translation process.
High-level code. High-level language programs require both similar and additional pre-processing operations to those performed on the assembly code. First, printed strings can be removed from the target programs since there is no solution for the model to predict them. Since we know that each program generates its own bad and good versions, one of the fundamental operations to perform is splitting them. For this purpose, we leveraged pre-compiler directives.
The next pre-processing step is comment removal. Similarly to low-level code pre-processing, we removed the main function and the inclusions of libraries and namespaces. Again, we replaced function names with standard names (e.g., funct0) to let the model focus exclusively on the algorithm of the actual functions.
Tokenization. After the pre-processing phase, both the high- and low-level code are represented as unstructured text, hence the need to tokenize it. The first stage of this process is a pre-tokenization to split the words of the training data according to the blanks. We adopted the Byte Pair Encoding (BPE) technique, typically preferred for tasks such as NMT (hosseini2022beyond, ). Then, we performed toss reduction, i.e., removal of programs that exceed a specific length in terms of the number of tokens: we cut off those programs whose size is above the 95th percentile ($1,479$ tokens for the source language, $338$ for the target) and below the 5th percentile ($20$ for the source language, $9$ for the target). As a result of this operation, the number of elements decreased from $83,520$ to $75,149$.
Table 3. Dataset pre-processing
Statistic Compiled Dataset Pre-processed Dataset
C/C++ ASM C/C++ ASM
Function Count $83,520$ $83,520$ $75,149$ $75,149$
CWE Count $89$ $89$ $89$ $89$
Avg. Function Length $612$ $1,821$ $87$ $158$
Max Function Length $1,148$ $17,680$ $338$ $1,479$
@@ 3.4. Neural Decompilation
Neural decompilation is an emerging approach to translating binary code into source code. The main advantages of this approach are the following:
  • The decompiled program control-flow structure traces the structure of the original high-level program, making the results human-readable. This happens because the transformer learns the typical structure of source programs.
  • Neural decompilation is a data-driven approach, hence it can be trained on new target languages or architectures with new samples set without depending on specific and hard-to-use toolchains.
Neural decompilation leverages Neural Machine Translation (NMT), a machine learning approach to translating a given sequence of elements into another sequence through the use of Large Language Models (LLMs), deep learning algorithms able to perform a set of language processing tasks such as translation, classification or generation. In light of the complexity of the tasks, LLMs generally require large accurate datasets to be trained. The core of the LLMs is the Transformer neural network, a deep learning architecture consisting of an encoder-decoder structure based on the multi-headed attention mechanism (vaswani2017attention, ), often used for translation from one language to another. In our case, we leverage NMT to translate machine code to high-level code. The main difference between the two uses is the number of tokens between the source and the translated text. There is a wide gap in the number of tokens between the machine and high-level code. Another difference is that, although natural language follows grammatical rules, a programming language is more structured by nature. The fine-tuned models for the translation task are the following:
  • CodeBERT (feng2020codebert, ): is a large multi-layer bidirectional Transformer with an encoder-only architecture pre-trained on millions of lines of code across six different programming languages created by Microsoft. For the translation task, we combined it with a decoder to have an encoder-decoder architecture suitable for code generation (wang2021codet5, ).
  • CodeT5+ (wang2023codet5, ): is a large multi-layer bidirectional Transformer with an encoder-decoder architecture, developed by Salesforce and pre-trained on diverse tasks including causal language modeling, contrastive learning, and text-code matching. We utilize the variant with model size $220M$, trained from scratch following T5’s architecture (wang2021codet5, ) and initialize it with a checkpoint further pre-trained on Python.
  • fairseq (fairseq, ): is a sequence modeling toolkit written in PyTorch that allows researchers and developers to train custom models for text translation and generation, created by Meta. This model is not pre-trained and supports encoder-decoder architectures for a wide variety of tasks. We selected this model on the basis of previous work (hosseini2022beyond, ).
@@ 3.5. Vulnerability Prediction
This stage represents the core of the vulnerability prediction task. The goal is to identify whether a target program is vulnerable, along with the type of vulnerability affecting it. We addressed these problems as classification problems, following the direction of several studies about vulnerability prediction (zheng2019recurrent, ; schaad2023deep, ; pewny2015cross, ). Specifically, we faced the following classification tasks:
@@
  • Multi-class classification: aims to identify which vulnerability affects the target program. Vulnerability types are the classes to be predicted by the classification models. Such vulnerabilities are mapped to the CWE (CWE, ), a standard taxonomy defined by MITRE (Mitre, ).
Therefore, we train several models on a training set consisting of original C/C++ programs, i.e., non-decompiled, and then test them on a test set consisting of neurally decompiled C/C++ programs. We make sure that overfitting does not occur since none of the samples in the original C/C++ training set are contained in the decompiled test set. Looking at previous studies on this topic (chen2023diversevul, ; schaad2023deep, ; zheng2019recurrent, ), we selected models that can handle the complexity of vulnerability prediction, including:
  • CodeBERT (feng2020codebert, ): unlike its use for the neural decompilation task, we used it as an encoder-only architecture for vulnerability prediction.
  • CodeT5+ (wang2023codet5, ): described in Section 3.4.
  • CodeGPT (huggingface2020codegpt, ): a Transformer based on the GPT-2 model family. CodeGPT provides a decoder-only architectural configuration. It was primarily designed for code generation and optimized for Python, but can also be adapted to generic code classification tasks.
  • Long Short-Term Memory (LSTM): a recurrent neural network with retroactive connections that provides short-term memory. This makes it fit for processing long input data sequences, as in the case of code classification.
  • Simple Recurrent Neural Networks (SRNN): it consists of a neural network with interconnections between units through cycles. This enables determining a temporal dynamicity, dependent on the information received at the previous times.
  • Gated Recurrent Unit (GRU): consists of a recurrent neural network that introduces a gating mechanism, i.e., a mechanism to decide which information is forwarded to the output.
@@
Proposed approach CodeBERT 75,000 338 31% 39% 43% 56%
CodeT5+ 75,000 338 48% 49% 55% 70%
fairseq 75,000 338 58% 59% 64% 77%
*: results obtained on a random test set of Juliet dataset.
°: trained with different data than Juliet.
@@ 4. Experimental Analysis
4.1. Experimental Setup
The analysis was performed on a virtual instance of Google Cloud running on Debian GNU/Linux OS version 10, equipped with an Intel Haswell CPU, 16 GB RAM, a 250 GB disk, and a Tesla V100-SXM2-16GB GPU.
Neural Decompilation. The translation task was performed using fairseq, CodeBERT, and CodeT5+ (base: 220M) models. CodeBERT is trained by implementing an encoder-decoder framework where the encoder is initialized with the pre-trained CodeBERT weights, while the decoder is a transformer decoder comprising 6 stacked layers. The encoder leverages the RoBERTa architecture (liu2019roberta, ), with 12 attention heads, 768 hidden layers, 12 encoder layers, and 514 for the size of position embeddings.
CodeT5+ has an encoder-decoder architecture with 12 decoder layers, each with 12 attention heads and 768 hidden layers, and 512 for the size of position embeddings. We set learning rates of 5e-5, batch size of 8, beam size of 10, evaluation steps of 314, and training steps of 2005 for both CodeBERT and CodeT5+ models. Source and target length were set to 512, the maximum length of the input sequences in both CodeBERT and CodeT5+ models. We truncated all the assembly programs longer than the maximum supported length.
We derived basic fairseq hyperparameters configuration for model training from the work of Hosseini et al. (hosseini2022beyond, ). We chose a different optimizer, namely the nag optimizer. We performed 10-fold cross-validation to evaluate the decompilation models using the whole dataset and have a more robust and reliable evaluation. Specifically, we trained and evaluated the models ten times, each using one of the ten folds as the test set and the remaining nine as the training set. At the end of the ten iterations, performance is averaged to obtain a more reliable estimate of the model’s predictive capabilities. Vulnerability Prediction. For the neural network-based models, namely LSTM, SRNN, and GRU, we used 4 bidirectional layers with a number of units equal to the maximum sequence length set during pre-processing and one dense layer with a number of layers equal to the number of classes for all three networks. For training, we employed the Categorical Cross-entropy metric and the Adam optimizer, with an initial learning rate of $41e-$ 4. We configured $100$ epochs, adding callbacks to improve and speed up learning. We set the patience value for Early Stopping to 5 epochs, hence the training was stopped if the loss did not grow within 5 epochs. Each training lasted between $30$ and $40$ epochs. We chose the Word2Vec Gensim library for pre-processing, i.e., from tokens to embeddings. We implemented input padding to prevent the model from basing its choices on the programs length.
On the other hand, for the transformer-based models, we set the batch size per device to 4 for both training and evaluation. The learning rate was set to $52e-$ 5, a low value to ensure gradual updates, employing a linear learning rate scheduler to progressively reduce the learning rate throughout training. The weight decay was set to 0.01 to regularize the model and prevent overfitting. The Adam optimizer was used with an epsilon of $81e-$ 8 to ensure numerical stability. We applied gradient accumulation in two steps, which would mean the gradients were accumulated in two iterations before updating the model. The models were trained for 3 epochs. We utilized the tokenizer models provided by the LLMs themselves for proper input processing: T5Tokenizer for T5, GPT2Tokenizer for CodeGPT, and RobertaTokenizer for CodeBERT.

### v2

@@ Abstract.
Vulnerability prediction is valuable in identifying security issues efficiently, even though it requires the source code of the target software system, which is a restrictive hypothesis. This paper presents an experimental study to predict vulnerabilities in binary code without source code or complex representations of the binary, leveraging the pivotal idea of decompiling the binary file through neural decompilation and predicting vulnerabilities through deep learning on the decompiled source code. The results outperform the state-of-the-art in both neural decompilation and vulnerability prediction, showing that it is possible to identify vulnerable programs with this approach concerning bi-class (vulnerable/non-vulnerable) and multi-class (type of vulnerability) analysis.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXbooktitle: The 18th European Workshop on Systems Security (EuroSec ’25), March 31, 2025, Rotterdam, Netherlandsisbn: 978-1-4503-XXXX-X/18/06journalyear: 2025conference: The 18th European Workshop on Systems Security; March 30–April 3, 2025; Rotterdam, Netherlandsbooktitle: The 18th European Workshop on Systems Security (EuroSec ’25), March 30–April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3722041.3723097isbn: 979-8-4007-1563-1/25/03ccs: Security and privacy Software security engineeringccs: Security and privacy Vulnerability management
@@ 1. Introduction
Neural decompilation is an emerging method that leverages modern deep learning techniques, such as transformers and self-attention, to translate binary code back into source code. One of the key benefits of this approach is that the control flow structure of the decompiled program resembles the original source code, in a closer way than disassemblers and traditional decompiler tools, such as Ghidra (Ghidra, ). Neural decompilation is a data-driven approach, where the neural network learns the typical patterns of source code structures. Gaining access to the source code offers several advantages, such as enabling vulnerability detection, malware recognition, re-engineering of legacy systems, and facilitating static analysis. Moreover, binary decompilation enables vulnerability prediction in “black-box” vulnerability assessments (e.g., proprietary software), where the assessor attacks the system from the same perspective as the attacker, without source code.
@@
In this work, we explore the potential to identify software vulnerabilities in binary files via neural decompilation (Figure 1), without need to access the source code of the target software. The results of this study show that:
@@
The remainder of the paper is structured as follows. In Section 2, we elaborate on background concepts about the target problem. In Section 3, we present the methodology for vulnerability prediction via neural decompilation. In Section 4, we illustrate the experimental analysis for the decompilation and vulnerability prediction tasks. In Section 5, we examine related work. Section 6 concludes the paper.
@@ 2. Background
The main complexities of this work lie in the decompilation process, i.e., the transformation of executable code into source code. This issue stems from the semantic gap between high- and low-level code: machine code does not provide function-level abstraction since function calls are implemented through call labels and jumps that allow the execution of loops and conditional constructs. This asymmetry causes the loss of the control flow of the original program, which is quite challenging to re-construct from the binary. Several decompilers are available in this field (e.g., Ghidra (Ghidra, ), RetDec (RetDec, )), achieving reasonable levels of accuracy; however, they come with high costs in terms of knowledge of the specific languages, generation of control flow graphs, and data flow graphs.
Another complexity in dealing with the translation and vulnerability prediction tasks lies in finding an appropriate dataset. The dataset should meet the following requirements:
@@
ReVeal (arewethereyet, ) 18,1691816918,16918 , 169 No N/A
Devign (zhou2019devign, ) 26,0372603726,03726 , 037 No N/A
Juliet (Juliet, ) 128,198128198128,198128 , 198 Yes Yes
BigVul (bigvul, ) 264,919264919264,919264 , 919 No Only for 8,783 samples
@@
As shown in Table 1, among the C/C++ datasets for vulnerability prediction available in the literature (chen2023diversevul, ; bigvul, ; arewethereyet, ; zhou2019devign, ), the only dataset suitable for our study is Juliet C/C++ 1.3 (Juliet, ), a synthetic collection of C/C++ programs from the Software Assurance Reference Dataset (SARD) (SARD, ), developed by NIST (NIST, ). The Juliet test suite was first released in 2010, gaining increasing popularity for benchmarking static analysis tools and vulnerability prediction tasks (zheng2019recurrent, ; schaad2023deep, ; li2021vuldeelocator, ; li2018vuldeepecker, ; jain2023code, ; gao2018comprehensive, ) since it includes both vulnerable and non-vulnerable programs. The dataset provides 64,0996409964,09964 , 099 compilable test cases for 118 CWEs. For each CWE, a test case comprises a program vulnerable to the target CWE and a non-vulnerable one.
@@ 3. Methodology
  • To avoid translating binary code to high-level code directly, we disassembled the binary into assembly code. Working directly with binary code would prevent information pre-processing. Consequently, the input for the decompilation model will be assembly code, further pre-processed in the next steps.
@@
Our experimental approach revolves around five phases, as shown in Figure 2. The first phase is Binary Generation to compile the C/C++ programs under analysis. We adopted the automated building approach by Richardson et al. (JulietCompilation), configuring the gcc/g++ compiler (depending on the program’s language) with the -O0 optimization level. Due to compilation issues, the dataset was reduced from 128,198 programs to 83,520. The second one is Disassembly, which lifts the binary code to assembly code via objdump (objdump, ). The format of the executables is elf64-x86-64. The third phase is Pre-processing of the low- and high-level code. In the low-level code, we removed all sections except .text since we are only interested in the algorithmic logic. We removed library functions defined within the same section, operating system directives, stack initialization, and the main function, which is constant in every program, making its contribution to learning negligible. In the high-level code, we removed the main function, library inclusions, namespaces, comments, and printed strings since there is no solution for the model to predict them. From both low- and high-level code, we replaced function names with standard names (e.g., funct0) to let the model focus exclusively on the algorithm of the actual functions. Finally, we applied toss reduction on tokenized code via BPE, removing programs above the 95th percentile and below the 5th percentile. This reduced the dataset from 83,520 to 75,149 elements. The fourth and fifth phases are Neural Decompilation and Vulnerability Prediction.
@@ 3.1. Neural Decompilation
Neural decompilation is an emerging approach to translating binary code into source code. The main advantages of this approach are two. The first is that the decompiled program control-flow structure traces that of the original high-level program, making the results human-readable. This happens because the transformer learns the typical structure of source programs. The second advantage is that neural decompilation is a data-driven approach, hence it can be trained on new target languages or architectures with new data without depending on specific and hard-to-use toolchains. Neural decompilation leverages Neural Machine Translation (NMT), a machine learning approach to translating a given sequence of elements into another sequence through the use of Large Language Models (LLMs), deep learning algorithms able to perform a set of language processing tasks such as translation, classification or generation. In light of the complexity of the tasks, LLMs generally require large accurate datasets to be trained. The core of the LLMs is the Transformer neural network, a deep learning architecture consisting of an encoder-decoder structure based on the multi-headed attention mechanism (vaswani2017attention, ), often used for translation from one language to another. In our case, we leverage NMT to translate machine code to high-level code. The fine-tuned models we used for the translation task are CodeBERT (feng2020codebert, ), CodeT5+ (wang2023codet5, ) and fairseq (fairseq, ). For CodeBERT, we combined it with a decoder to have an encoder-decoder architecture suitable for code generation (wang2021codet5, ).
@@ 3.2. Vulnerability Prediction
This stage represents the core of the vulnerability prediction task. The goal is to identify whether a target program is vulnerable, along with the type of vulnerability affecting it. We addressed these problems as classification problems, following the direction of several studies in the field (zheng2019recurrent, ; schaad2023deep, ; pewny2015cross, ). Specifically, we faced the following classification tasks:
@@
  • Multi-class classification: aims to identify which vulnerability affects the program. Vulnerability types are the classes that the classification models predict. Such vulnerabilities are mapped to the CWE (CWE, ).
Therefore, we train several models on a training set consisting of original C/C++ programs, i.e., non-decompiled, and then test them on a test set consisting of neurally decompiled C/C++ programs. Looking at previous studies on this topic (chen2023diversevul, ; schaad2023deep, ; zheng2019recurrent, ), we selected models that can handle the complexity of vulnerability prediction, including CodeBERT (feng2020codebert, ), CodeT5+ (wang2023codet5, ), CodeGPT (huggingface2020codegpt, ), Simple Recurrent Neural Networks (SRNN), Long Short-Term Memory (LSTM) and Gated Recurrent Unit (GRU).
@@
Proposed approach CodeBERT 75,149 338 31% 39% 43% 56%
CodeT5+ 75,149 338 48% 49% 55% 70%
fairseq 75,149 338 58% 59% 64% 77%
*: results obtained on a random test set of Juliet dataset.                  °: trained with different data than Juliet.